OpenStudy (anonymous):

find the first 30 terms of the sequence an=4n+1??

OpenStudy (anonymous):

I need this answer ASAP!!!

OpenStudy (anonymous):

ok, well just in case you get this, you can find the first term by substituting 1 for n in the above statement. so, a1=4(1)+1=5, a2=4(2)+1=9.... and so on to a30. I hope this helps. take care